If you invest the $10,000 you receive at graduation in a mutual fund which averages a 12 % annual return, how much will have forty years later?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=10,000*(1.12)^40

=10,000*93.0509704

=$930509.70(Approx).